Teacher A has 190 students in her mathematics class. the scores in the final exam are normally distributed with a mean of 74.4 and standard deviation of 8.7.how many students in the class can be expected to receive a score between 84 and 90?
Answer by Boreal(15235)

z=(x-mean)/sd
lower value for z=(84-74.4)/8.7
z lower=+1.10
z upper=(90-74.4)/8.7=15.6/8.7=+1.79
probability of z being 1.10 and 1.79 is 0.0989
Multiply that by 190 (I didn't round until the end) and the answer is 19. Rounded, it would still be 19.